OpenStudy (anonymous):

which one is a prime number? 35, 69, 89, 90

OpenStudy (welshfella):

try and factor each number in turn if you cant find any its a prime

OpenStudy (welshfella):

eg 35 = 5*7 - not prime

OpenStudy (studygurl14):

A prime number cannot be divisible by any number except for 1 and itself.

OpenStudy (thedj4jc):

A prime number is a number that can only be divided by one and itself, find which of the numbers only has 1 and itselfs that it can be divided by

OpenStudy (welshfella):

handy hint: if you add up the digits of a number until you get 1 digit and its 3,6 or 9 then that number is divisible by 3

OpenStudy (welshfella):

foe example take 345 : 3+4+5 = 12 1 +2 = 3 so its divisible by 3 try that on the last 3 numbers in your question

OpenStudy (welshfella):

6 + 9 = ?

OpenStudy (anonymous):

15

OpenStudy (welshfella):

and 1 + 5 = ?

OpenStudy (anonymous):

6

OpenStudy (welshfella):

- so 69 can be divided by so its not prime

OpenStudy (welshfella):

* divided by 3

OpenStudy (january123):

35 = 5*7, 69 = 3*23, 90 = 2*3^2*5 and 10 = 2*5.

OpenStudy (january123):

35 = 5 * 7 69 = 3 * 23 90 = 2 * 45 10 = 2 * 5 ....and 89 is NOT divisible by 2, 3, 5, or 7 (primes less than the square root of 89)

OpenStudy (anonymous):

The prime one is the one that it self and 1 can only go in to it for example 36 would not be prime because 36 and 1 and 6, 2,3,4 and 9 can go into it

OpenStudy (anonymous):

so is it 89?

OpenStudy (anonymous):

yes very good